OpenStudy (gtxmuqsit):

a refrigerator has to transfer an average of 263 j of heat per secound from -10 to 25 centgrade. calculate the average power consumed. asumed that the refrigerator is ideal a.25w b.30w c.35w d.40w e.45w @AllTehMaffs

OpenStudy (anonymous):

that's the question @sarah786 posted, right? I can't figure out how to do this one for some reason.

OpenStudy (gtxmuqsit):

yeah @AllTehMaffs it was in my test I was unable to solve it

OpenStudy (anonymous):

Do you have any idea @Lessis ?

OpenStudy (anonymous):

It has something to do with reversed Carnot cycles First law of thermo is \[Q_h=Q_c+W\] and we want \[P_{avg} = \frac{W}{t}\] But it gives you a power in the first place, right? 263J/s Just can't figure out how to relate that to anything :/

OpenStudy (gtxmuqsit):

@AllTehMaffs it's tough

OpenStudy (anonymous):

it's driving me crazy 0_o

OpenStudy (anonymous):

I'm not sure about what you're supossed to do with the given temperatures. Also, an ideal refrigerator is 100% efficient. Right?

OpenStudy (anonymous):

yeah \[COP = 1 = \frac{T_c}{T_h-T_c}=\frac{Q_c}{W}\]

OpenStudy (anonymous):

\[Efficiency = \frac{ Q_{1} }{ Q_{2}-Q_{1} }=\frac{ Q_{1} }{ W } \]

OpenStudy (anonymous):

for a fridge, yah

OpenStudy (anonymous):

\[ \frac{Q_h}{Q_c} \propto \frac{T_h}{T_c}\]

OpenStudy (anonymous):

\[\frac{1}{\frac{Q_h}{Q_c} - 1} \propto \frac{1}{\frac{T_h}{T_c}-1} = \frac{T_c}{T_h-T_c}\]

OpenStudy (anonymous):

err, \[ \frac{Q_h}{Q_c}≥\frac{T_h}{T_c}\]

OpenStudy (anonymous):

Second law of thermo - entropy can increase but not decrease, so in a heat engine, the entropy it expels has to be at least a much as the entropy it absorbs entropy = \[S=\frac{Q}{T}\] \[S_c=\frac{Q_c}{T_c} \\ \ \\ S_h=\frac{Q_h}{T_h}\] \[S_c≥S_h\] \[\frac{Q_c}{T_c}≥\frac{Q_h}{T_h}\] \[\frac{Q_c}{Q_h}≥\frac{T_c}{T_h}\]

OpenStudy (anonymous):

but still dunno what to do with that freaking 263J/s :P

OpenStudy (anonymous):

Huh. Well, the answer is 35 W. I guess. xD

OpenStudy (anonymous):

I THINK THIS IS WHAT YOU HAVE TO DO. \[COP = \frac{ T_{c} }{ T_{h} - T_{c} } \approx 7.51\] \[COP = \frac{ Q_{c} }{ W }\] \[\frac{ Q_{c} }{ W } = \frac{ Q_{c}/t }{ P } = COP\] \[P = \frac{ Q_{c}/t }{ COP } \approx 35 W\]

OpenStudy (anonymous):

WINNER!!!

OpenStudy (anonymous):

Oh. And Qc/t is the head rejected per second, which is 263 J/s.

OpenStudy (anonymous):

ooooohhhh

OpenStudy (anonymous):

@sarah786 an answer!!

OpenStudy (anonymous):

I really just divided 263/7.5 to see what happened though. xD And then I just deduced the equation from there.

OpenStudy (anonymous):

guess and check is the best way to math ^_^

OpenStudy (anonymous):

Good thing this was a multiple choice question, huh? D:

OpenStudy (anonymous):

For reals - the one benefit of them :P

OpenStudy (anonymous):

Thank you so much @Lessis @AllTehMaffs

OpenStudy (gtxmuqsit):

@AllTehMaffs @Lessis thanks :P